# LatestEval: Addressing Data Contamination in Language Model Evaluation through Dynamic and Time-Sensitive Test Construction

### 26 Dec 2023 v2

@@ Introduction
Recent years have seen the ubiquity of pretrained language models in natural language processing (NLP) due to their strong performance and generalisation capability. These models are usually pre-trained on super large internet-crawled corpora. However, many widely used benchmarks are also largely constructed from web resources (Hendrycks et al. 2021), which are very likely to be unintentionally included in the pretraining stage. This leads to a major emerging issue called data contamination.
Recent analysis has revealed that data contamination is widespread in model evaluations (OpenAI 2023; Anil et al. 2023), which greatly undermines the credibility of evaluation results (Marie 2023; Narayanan and Kapoor 2023) and prevents fair comparisons between models (Dickson 2023). Moreover, the massive scale of training data makes decontaminating existing benchmarks extremely difficult (Kreutzer et al. 2022). For many closed models, training data is considered as trade secret and thus confidential, eliminating any possibility for the community to address contamination by decontaminating benchmarks. One potential solution to avoid contaminated evaluation is to create new test data constantly or use human evaluation (Liu, Zhang, and Liang 2023; Jacovi et al. 2023), just like how examination for human works. However, this is extremely inefficient and costly, requiring huge human efforts periodically.
@@
What is data contamination?   Data contamination refers to the phenomenon that examples from the evaluation set are also found in the training data. This might lead to the evaluation failing to accurately reflect models’ capabilities, as models can cheat by memorising instead of learning to generalise. There are two primary types of data contamination (Dodge et al. 2021): (i) input contamination refers to only the input appearing in the pretraining corpus, and (ii) input-and-label contamination is when both inputs and their labels are present. The latter is generally more problematic, as models can directly memorise input-output pairs. However, the former can still cause issues as models may gain an advantage by learning from the input, especially for assessing few-shot and zero-shot learning capabilities.
@@
How common is data contamination?   Data contamination appears to be quite widespread across commonly used NLP benchmark datasets based on findings from recent studies. Dodge et al. (2021) revealed exact match contamination rates ranging from under 2% to over 50% on various GLUE benchmarks when compared to the C4 pretraining data. The GPT-3 study (Brown et al. 2020) found over 90% of examples in Quac, SQuADv2, and DROP were flagged as contaminated. FLAN (Wei et al. 2021) evaluations identified 7 out of 26 datasets exhibiting a serious contamination ratio of 50% and over. LLaMA 2 (Touvron et al. 2023) reported over 16% of MMLU examples are contaminated and about 11% are seriously contaminated (more than 80% token leakage). GPT-4 (OpenAI 2023) uses academic exams instead of NLP benchmarks for model evaluation. While 4 out of 34 exams are found to have zero contamination (e.g., Leetcode and Bar Exam), 9 out of 34 showed over 20% of instances are marked as dirty examples. Li (2023b) identified varying level of data contamination ranging from 1-46% in six popular English and multilingual multi-choice QA benchmarks.
@@
How to identify data contamination?   Dodge et al. (2021) takes a straightforward approach to detect exact matches between test set examples and the pretraining data after normalising for capitalisation and punctuation. The exact match here means the entire input of an evaluation text is found in the training data. The GPT-3 paper (Brown et al. 2020) uses n-gram overlap to identify contamination, treating any examples with 13-gram co-occurrence in both test sets and training data as dirty examples. PaLM (Chowdhery et al. 2022) considers a sample to be contaminated if 70% of its 8-grams can be found at least once in the training data. LLaMA-2 matches on tokenized prompts and takes a bottom-up, token-level approach to identify contamination. Overall, existing approaches usually use substring matching between evaluation examples and training data to identify data contamination. However, if we have no access to the training data, which is often the case for most recent closed models, it is extremely difficult to reveal contamination by observing the models themselves. Some pioneering studies (Carlini et al. 2021, 2022; Li 2023a) propose an approach to check whether a model has memorised an example by asking it to reconstruct the example verbatim, which can potentially used to detect evaluation data leakage. Li (2023b) introduced an open source pipeline to identify potential data contamination with search engine, and audit model performance separately on the clean and contaminated test set.
@@
To what extent does data contamination affect model evaluation?   While contaminated data can potentially inflate scores, models do not necessarily perform worse on clean subsets or better on dirty subsets across all datasets. The degree of impact likely depends on factors like the dataset characteristics, model scale, and nature of the pre-training data. For instance, GPT-3 (Brown et al. 2020) showed a small 1-2% performance drop on clean subsets for PIQA and ReCoRD, compared to a significant 6% drop on clean set of SQuAD as 94% of its test examples were contaminated. The LLaMA model (Touvron et al. 2023) did not show significant gaps between clean and dirty subset performance. On HellaSwag, LLaMA’s 70B model showed a 15.3 point gap between clean (63.5) and dirty (78.8) subsets. Detecting and accounting for data contamination remains an active area of research, as there is no consensus yet on best methodologies and acceptable contamination levels.
@@
First, we conduct a contamination test on common reading comprehension benchmarks to measure the extent that models memorise their test instances during pretraining. Although most existing analyses identify contamination by computing the overlapping between training and test sets, we cannot use the same approach as the pretraining data of current language models are mostly not publicly available. Therefore, we propose a novel method to quantify test contamination by observing whether models exhibit memorisation behaviour on test instances, which can serve as a strong signal of test contamination. Carlini et al. (2021, 2022) have defined the “memorisation” of models that a sequence is considered as memorised if the model has considerably smaller perplexity on that sequence. The idea is that sequences leak in the training data will tend to have lower perplexity (i.e., higher likelihood) than sequence models never seen before. This is well demonstrated in Figure 2, where we compare the perplexity of two Wikipedia subsets: the first is Wikitext (Merity et al. 2016) which is widely used in language model pretraining, the second is NewWiki, latest Wikipedia texts that created after April 2023, after all tested models were released. We find that Wikipedia content learned during pretraining has a considerably lower perplexity than recently created Wikipedia content.
@@
Now we compute the perplexity of the validation set of the three existing reading comprehension benchmarks SQuAD v2 (Rajpurkar, Jia, and Liang 2018), BoolQ (Clark et al. 2019), QuAC (Choi et al. 2018) and compare them to LatestEval. We involve five foundation models here: opt-350m, opt-1.3b (Zhang et al. 2022), Llama-7b, Llama-13b, and Llama-30b (Touvron et al. 2023). OPT models are run with full precision and Llama models are run with fp16. We make sure the input texts of each benchmark have the same length so that we can obtain a fair comparison.

### 01 Mar 2024 v3

@@ Introduction
Recent years have seen the ubiquity of pretrained language models in natural language processing (NLP) due to their strong performance and generalisation capability. These models are usually pre-trained on super large internet-crawled corpora. However, many widely used benchmarks are also largely constructed from web resources (Hendrycks et al. 2020), which are very likely to be unintentionally included in the pretraining stage. This leads to a major emerging issue called data contamination.
Recent analysis has revealed that data contamination is widespread in model evaluations (Achiam et al. 2023; Sainz et al. 2023), which greatly undermines the credibility of evaluation results (Marie 2023; Li, Guerin, and Lin 2023) and prevents fair comparisons between models (Dickson 2023). Moreover, the massive scale of training data makes decontaminating existing benchmarks extremely difficult (Kreutzer et al. 2022). For many closed models, training data is considered as trade secret and thus confidential, eliminating any possibility for the community to address contamination by decontaminating benchmarks. One potential solution to avoid contaminated evaluation is to create new test data constantly or use human evaluation (Liu, Zhang, and Liang 2023; Jacovi et al. 2023), just like how examination for human works. However, this is extremely inefficient and costly, requiring huge human efforts periodically.
@@
What is data contamination?   Data contamination refers to the phenomenon that examples from the evaluation set are also found in the training data. This might lead to the evaluation failing to accurately reflect models’ capabilities, as models can cheat by memorising instead of learning to generalise. There are two primary types of data contamination (Dodge et al. 2021): (i) input-only contamination refers to only the input appearing in the pretraining corpus, and (ii) input-and-label contamination is when both inputs and their labels are present. The latter is generally more problematic, as models can directly memorise input-output pairs. However, the former can still cause issues as models may gain an advantage by learning extra information from the context (Li, Guerin, and Lin 2023).
@@
How common is data contamination?   Data contamination appears to be quite widespread across commonly used NLP benchmark datasets based on findings from recent studies. Dodge et al. (2021) revealed exact match contamination rates ranging from under 2% to over 50% on various GLUE benchmarks when compared to the C4 pretraining data. The GPT-3 study (Brown et al. 2020) found over 90% of examples in Quac, SQuADv2, and DROP were flagged as contaminated. FLAN (Wei et al. 2021) evaluations identified 7 out of 26 datasets exhibiting a serious contamination ratio of 50% and over. LLaMA 2 (Touvron et al. 2023) reported over 16% of MMLU examples are contaminated and about 11% are seriously contaminated (more than 80% token leakage). GPT-4 (Achiam et al. 2023) uses academic exams instead of NLP benchmarks for model evaluation. While 4 out of 34 exams are found to have zero contamination (e.g., Leetcode and Bar Exam), 9 out of 34 showed over 20% of instances are marked as dirty examples. Li, Guerin, and Lin (2023) illustrate varying level of contamination ranging from 2% to 45% in MMLU, C-Eval (Huang et al. 2023), Hellaswag (Zellers et al. 2019) and other popular QA benchmarks.
@@
How to identify data contamination?   Dodge et al. (2021) takes a straightforward approach to detect exact matches between test set examples and the pretraining data after normalising for capitalisation and punctuation. The exact match here means the entire input of an evaluation text is found in the training data. The GPT-3 paper (Brown et al. 2020) uses n-gram overlap to identify contamination, treating any examples with 13-gram co-occurrence in both test sets and training data as dirty examples. LLaMA-2 matches on tokenized prompts and takes a bottom-up, token-level approach to identify contamination. Overall, existing approaches usually use substring matching between evaluation examples and training data to identify data contamination. However, if we have no access to the training data, which is often the case for most recent closed models, it is extremely difficult to reveal contamination by observing the models themselves. Pioneering studies propose to identify benchmark data contamination by using search engine (Li, Guerin, and Lin 2023), measuring perplexity of test examples (Li 2023), or asking models to reconstruct test examples verbatim (Carlini et al. 2021, 2022).
@@
To what extent does data contamination affect model evaluation?   While contaminated data can potentially inflate scores, models do not necessarily perform worse on clean subsets or better on dirty subsets across all datasets. The degree of impact likely depends on factors like the dataset characteristics, model scale, and nature of the pre-training data. For instance, GPT-3 (Brown et al. 2020) showed a small 1-2% performance drop on clean subsets for PIQA and ReCoRD, compared to a significant 6% drop on clean set of SQuAD as 94% of its test examples were contaminated. The LLaMA model (Touvron et al. 2023) did not show significant gaps between clean and dirty subset performance. On HellaSwag, LLaMA’s 70B model showed a 15.3 point gap between clean (63.5) and dirty (78.8) subsets. Li, Guerin, and Lin (2023) reveal contamination can inflate benchmark score even if the contamination does not give away the answer (i.e., input-only contamination). It also find larger models obtain more advantages from data contamination due to their more powerful memorisation.
@@
First, we conduct a contamination test on common reading comprehension benchmarks to measure the extent that models memorise their test instances during pretraining. Although most existing analyses identify contamination by computing the overlapping between training and test sets, we cannot use the same approach as the pretraining data of current language models are mostly not publicly available. Therefore, we propose a novel method to quantify test contamination by observing whether models exhibit memorisation behaviour on test instances, which can serve as a strong signal of test contamination. Carlini et al. (2021, 2022) have defined the “memorisation” of models that a sequence is considered as memorised if the model has considerably smaller perplexity on that sequence. The idea is that sequences leak in the training data will tend to have lower perplexity (i.e., higher likelihood) than sequence models never seen before. This is well demonstrated in Figure 2, where we compare the perplexity of two Wikipedia subsets: the first is Wikitext (Merity et al. 2016) which is widely used in language model pretraining, the second is NewWiki, latest Wikipedia texts that created after April 2023, after all tested models were released. We find that Wikipedia content learned during pretraining has a much lower perplexity than recently created Wikipedia content.
@@
Now we compute the perplexity of the validation set on the three existing reading comprehension benchmarks SQuADv2 (Rajpurkar, Jia, and Liang 2018), BoolQ (Clark et al. 2019), QuAC (Choi et al. 2018) and compare them to LatestEval. We involve five foundation models here: opt-350m, opt-1.3b (Zhang et al. 2022), Llama-7b, Llama-13b, and Llama-30b (Touvron et al. 2023). OPT models are run with fp32 and others are run with fp16. We make sure the input texts of each benchmark have the same length so that we can obtain a fair comparison.